In five years time, I will be three times as old as I was three years ago, how old am I now?
a) Write an equation using variables to represent the above statement.


b) Solve your equation to discover the age.

Answer by chibisan(131)

a)
let your present age be x
after 5years ,
5+x = 3(x-3)

b)
5+9=3x-x
14=2x
7=x (your present age)